OpenStudy (anonymous):

Simplify only using positive exponents... (4^-1 + 2^-1)^2

OpenStudy (anonymous):

\[(\frac{1}{4}+\frac{1}{2})^2\] \[=(\frac{3}{4})^2=\frac{9}{16}\]

OpenStudy (anonymous):

you need to know that \[4^{-1}=\frac{1}{4}\] because \[b^{-n}=\frac{1}{b^n}\]